# radarODE-MTL: A Multi-Task Learning Framework with Eccentric Gradient Alignment for Robust Radar-Based ECG Reconstruction

### v1

@@ IV-B2 Indoor Scene Understanding
Figure 5: Illustration of different frameworks under the same abrupt RBM noise: (a) MMECG cannot provide reasonable recovery under the noise; (b) radarODE could provide the general ECG shape but with a misalignment issue; (c) radarODE-MTL generates the high-fidelity ECG recovery under the noise.
To conclude the above evaluations in terms of different tasks, the proposed EGA could successfully alleviate the gradient conflicts and magnitude dominance in MTL optimization, while the intrinsic task difficulty can be successfully estimated to guide the optimization direction by introducing eccentric vector $\boldsymbol{v}_{ecc}$. Compared with other methods, EGA achieves an outstanding result for the tasks with disparate difficulties and is also competitive in the common tasks, but the hyperparameter $T$ should be carefully selected with large $T$ evenly treating all the tasks and small $T$ highlighting the hard task in each training epoch.
Figure 6: Evaluations for long-term ECG recovery: (a) - (d) CDF plots of MDR, HR Error, RMSE and PCC, with corresponding improvements.
@@ IV-C Evaluations on the Long-term Recovered ECG
The outputs from Task $1-3$ can form the long-term ECG signal as introduced in Methodology and the general performance can be depicted in Figure 5, with all three frameworks successfully reconstructing the ECG signals without noises. However, Figure LABEL:sub@fig:mmecg_noise shows that MMECG cannot resist RBM noise and obtains bad RMSE/PCC as also reported in the benchmark paper [2]. In contrast, radarODE could preserve the general shapes of ECG pieces due to the constrain from ODE model [5], but the peaks deviate from the ground truth because the inaccurate PPI estimation and the reintroduction of noises in long-term reconstruction stage, degrading the RMSE/PCC as shown in Figure LABEL:sub@fig:ode_noise. Lastly, the proposed radarODE-MTL realizes the ECG reconstruction in an end-to-end manner without reintroducing the noises, and the recovered ECG is less corrupted by the noises with the best RMSE/PCC compared with other frameworks as shown in Figure LABEL:sub@fig:mtl_noise. The following part will thoroughly evaluate the recovered long-term ECG signals in terms of corrupt ECG reconstruction, coarse cardiac feature and fine-grained ECG feature reconstruction, respectively.
IV-C1 Corrupt ECG Reconstruction
The concept of MDR is used to statistically evaluate the corruptions in the recovered ECG signals due to noise distortion. The result is shown as the cumulative distribution function (CDF) of MDR in Figure LABEL:sub@fig:mdr_cdf with the median MDR as $1.7\%$, $0.13\%$ and $0.13\%$ for MMECG, radarODE and radarODE-MTL respectively, and $\Delta m\%$ across $91$ trails are both $14\%$. The reason for the similar performance of two ODE-based methods is that the misaligned ECG pieces with small deviations ($<150$ms) in radarODE will not be identified as ‘missed detected’, and hence the CDFs of MDR share a similar pattern and trend in Figure LABEL:sub@fig:mdr_cdf.
@@ IV-C1 Corrupt ECG Reconstruction
IV-C2 Coarse Cardiac Feature Reconstruction
All three frameworks evaluated in this paper are designed for fine-grained cardiac features reconstruction and should perform well on the coarse cardiac feature (i.e., heart rate (HR) monitoring). The result in Figure LABEL:sub@fig:hr_cdf coincides with the expectation with median HR error as $0.6$, $0.3$ and $0.3$ beats/min respectively, and $\Delta m\%$ for the ODE-based methods are $43\%$ and $48\%$. It is notable in Figure LABEL:sub@fig:hr_cdf that the performances of ODE-based methods are very similar at the beginning, while the radarODE tends to get more errors when the noise in the raw radar signal affects the R peaks recovery, because the calculation of HR is based on the R peak positions.
IV-C3 Fine-Grained Morphological Feature Reconstruction
The morphological feature is an essential fine-grained feature to describe the general similarity between the recovered and ground truth ECG signals, and the morphological accuracy can be evaluated by RMSE and PCC, with RMSE sensitive to the peak deviation and PCC focusing on the similarity of the general shape. The results are shown in Figure LABEL:sub@fig:rmse_cdf and LABEL:sub@fig:ocor_cdf as the CDF of RMSE/PCC across $91$ trails in the dataset, and three frameworks get the median RMSE/PCC as $0.125$mV/$82.1\%$, $0.098$mV/$90.1\%$ and $0.083$mV/$92.7\%$ respectively. As indicated by $\Delta m\%$, the improvements of RMSE ($28\%,33\%$) are larger than PCC ($18\%,21\%$) for radarODE and radarODE-MTL respectively, because the ODE model embedded in the decoder preserves the main features of ECG even under noises and contributes more on the peaks than on the shapes. In addition, radarODE-MTL further improves the results by aligning the ECG pieces with the predicted anchors, avoiding the misalignment issue in radarODE.
IV-C4 Fine-Grained ECG Peaks Reconstruction
@@ IV-D Noise Robustness Test
In this work, $10$ trails (No. $75-84$) are selected for the noise robustness test by adding different types of synthesized noises with certain decibel (dB).
@@ IV-D1 Constant Noise
The constant noise normally affects the SNR of the signal, and can be simulated by adding Gaussian noise with different intensities as implemented in [6, 19, 40, 41, 42]. The baseline results for three frameworks are firstly obtained in terms of the RMSE, PCC, R-peak error and MDR as shown in Table III, and $\Delta m\%$ is calculated as $0\%$, $7.47\%$ and $10.64\%$ as indicated by the initial points in Figure 8. Then, the Gaussian noises with $6$ to $-3$dB are added into the raw radar signal without retraining the deep-learning framework, and the results are shown in Table III with the trends of performance degradation shown in Figure 8.
@@
SNR RMSE (mV) $\downarrow$ PCC $\uparrow$ Peak Error (ms) $\downarrow$ MDR $\downarrow$ $\Delta m\%^{1}\uparrow$
@@
Baseline 0.107 83.75% 9.45 4.52% 0.0%
${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}6$ dB 0.107 82.60% 9.76 4.37% -0.44%
${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}3$ dB 0.108 82.64% 9.85 4.84% -3.20%
${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}0$ dB 0.109 80.00% 11.80 4.92% -10.14%
$-1$ dB 0.114 78.55% 12.20 5.32% -15.02%
$-2$ dB 0.120 74.32% 14.64 5.59% -25.42%
$-3$ dB 0.127 62.45% 21.28 6.40% -52.82%
@@
Baseline 0.091 83.53% 9.08 4.03% 0.0%
${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}6$ dB 0.093 83.30% 9.12 4.36% -2.83%
${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}3$ dB 0.095 83.01% 9.01 4.70% -5.23%
${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}0$ dB 0.101 82.21% 9.89 5.86% -16.86%
$-1$ dB 0.116 79.66% 11.90 5.36% -24.17%
$-2$ dB 0.157 70.87% 13.95 6.19% -48.74%
$-3$ dB - - - - Failed2
@@
Baseline 0.089 85.03% 8.22 4.08% 0.0%
${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}6$ dB 0.088 85.31% 8.18 4.20% -0.31%
${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}3$ dB 0.089 84.29% 8.31 4.27% -1.87%
${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}0$ dB 0.091 83.77% 8.03 4.76% -4.58%
$-1$ dB 0.093 84.01% 8.10 5.10% -7.33%
$-2$ dB 0.093 84.51% 8.02 5.45% -9.18%
$-3$ dB 0.094 84.96% 8.19 6.02% -13.30%
@@
2. The ECG recovery fails if PCC$<60\%$, according to the empirical
    observation of the morphological ECG features.
A general observation of Table III is that all the frameworks perform well before 00dB with a similar degradation rate as in Figure LABEL:sub@fig:constant. Then, radarODE-MTL could still provide reasonable results with mild degradation after 00dB because the MTL paradigm split the ECG reconstruction task into several sub-tasks, and each task can either be constrained by prior knowledge or leverage the information from context data with less pollution. In contrast, radarODE could generate high-fidelity ECG pieces as claimed in [5] and gets the second best baseline result in Table III, but the design of PPI estimation stage does not consider the noise robustness. Therefore, the performance is heavily dropped to the worst in Figure LABEL:sub@fig:constant because of the bad results of Peak Error as shown in Table III. Lastly, the MMECG considers the ECG recovery as an arbitrary domain transformation problem without any constraints in the network design, and the performance also heavily degrades in Figure LABEL:sub@fig:constant because only meaningless results will be generated as shown previously in Figure LABEL:sub@fig:mmecg_noise.
@@ 2. The ECG recovery fails if PCC$<60\%$, according to the empirical
Duration RMSE (mV) $\downarrow$ PCC $\uparrow$ Peak Error (ms) $\downarrow$ MDR $\downarrow$ $\Delta m\%^{1}\uparrow$     RMSE (mV) $\downarrow$ PCC $\uparrow$ Peak Error (ms) $\downarrow$ MDR $\downarrow$ $\Delta m\%\uparrow$
@@ IV-D2 Abrupt Noise
In this part, the Gaussian noises with different intensities (00 and $-9$dB) are used to simulate mild body movement (e.g., during talking or writing) and extensive body movement (e.g., during torso movement) as suggested in the literature [43]. Only $20\%$ of the segments randomly selected from one trial are doped, and the duration of noise varies from $1$ to $3$ sec.
For mild body movement, the experimental results are shown in Table IV with the changes of $\Delta m\%$ shown in Figure LABEL:sub@fig:abrupt. Firstly, it is evident that the impact of $1$-sec abrupt noise is limited for all the frameworks, and the results for ODE-based methods are almost equivalent to the baselines. Secondly, $2$-sec noise starts to have a noticeable impact on MMECG, while the ODE-based methods could preserve the performance on the morphological features (RMSE/PCC) with small degradation on the Peak Error and MDR. Lastly, $3$-sec noise has distorted $3/4$ of the input radar segment, and the performances of MMECG and radarODE drop obviously as shown in Figure LABEL:sub@fig:abrupt, while radarODE-MTL only loses some points on $\text{MDR}=5.12\%$ as shown in Table IV.
In comparison, the extensive body movements with $1$ and $2$ sec have similar impacts with mild ones on ODE-based methods, because the ODE decoder could preserve the ECG shape even under strong noises, whereas the segments affected by noise cannot contribute to the recovery for MMECG as evident by the significant drop of PCC (from $84.05\%$ to $79.01\%$) as shown in Table IV. In addition, the $3$-sec noise destroys the ECG recovery for MMECG and radarODE with a significant degradation as shown in Figure LABEL:sub@fig:abrupt, whereas the radarODE-MTL only sacrifices certain RMSE and peak accuracy with the overall degradation dropping slightly from $-7.39\%$ to $-8.74\%$ as shown in Table IV.

### v2

@@ IV-B2 Indoor Scene Understanding
Statistical Analysis for EGA: To verify the significance of the EGA performance, the well-known T-test is performed with the null hypothesis as the performances of the compared methods are identical. The yielded $P$ values for all the methods are shown in Table I and II, and all the methods achieve $P<0.05$ except for using the equal weights, indicating that the obtained mean values are reliable with significance. In addition, it is worth noticing that the CIs of the loss balancing methods are larger than that of the gradient balancing methods for both tasks as shown in Table I and II, coinciding with the previous conclusion that the access of the gradients is beneficial to the stability of model training. Lastly, the proposed EGA not only achieves the best performance in ECG recovery, but the obtained results are also stable with small CIs, owing to the orthogonal projection that decreases the condition number of the gradient system [22].
To conclude the above evaluations in terms of different tasks, the proposed EGA could successfully alleviate the gradient conflicts and magnitude dominance in MTL optimization, while the intrinsic task difficulty can be successfully estimated to guide the optimization direction by introducing eccentric vector $\boldsymbol{v}_{ecc}$. Compared with other methods, EGA achieves an outstanding result for the tasks with disparate difficulties and is also competitive in the common tasks, but the hyperparameter $T$ should be carefully selected. During practice, the $T$ value can be adjusted until achieving optimum based on the fact that large $T$ evenly treats all the tasks and small $T$ enhances the hard task.
@@ IV-C Evaluations on the Long-term Recovered ECG
IV-C1 General Visualization for ECG Reconstruction
The outputs from Task $1-3$ can form the long-term ECG signal as depicted in Figure 6. All three frameworks successfully reconstruct the ECG signals from high-quality radar signals as shown in Figure LABEL:sub@fig:mmecg_goodLABEL:sub@fig:ode_good and LABEL:sub@fig:mtl_good, only with certain fluctuations in MMECG result and causing low RMSE/PCC/$R^{2}$.
In the presence of constant or abrupt noise, the signal SNR will decrease with the subtle features (e.g., $v_{2}$) ruined as shown in Figure LABEL:sub@fig:radar_moderate and LABEL:sub@fig:radar_abrupt. MMECG shows the least noise robustness and cannot resist abrupt noise as also reported in the benchmark paper [2] as shown in Figure LABEL:sub@fig:mmecg_moderate and LABEL:sub@fig:mmecg_abrupt, while radarODE achieves robust ECG recovery within each single cardiac cycle but shows obvious misalignment due to the PPI estimation error as shown in Figure LABEL:sub@fig:ode_moderate and LABEL:sub@fig:ode_abrupt. Lastly, the proposed radarODE-MTL realizes the ECG reconstruction in an end-to-end manner without reintroducing the noises, and the recovered ECG is less corrupted by the noises as shown in Figure LABEL:sub@fig:mtl_moderate and LABEL:sub@fig:mtl_abrupt.
Figure 7: Evaluations for long-term ECG recovery: (a) - (e) CDF plots of MDR, HR Error, RMSE, PCC and $R^{2}$, with corresponding improvements.
IV-C2 Corrupt ECG Reconstruction
The successful reconstruction in Figure LABEL:sub@fig:mtl_abrupt owes to the design of radarODE-MTL with deconstructed tasks for ECG recovery. Different from other frameworks with equal length of input and output, radarODE-MTL adopts a 4-sec segment to reconstruct the ECG piece for one cardiac cycle, and the radar signal from adjacent cardiac cycles (e.g., the zoomed part in Figure LABEL:sub@fig:radar_abrupt) also contributes to the recovery of the current ECG piece. In addition, if the input radar signal is fully destroyed by noise, radarODE-MTL may fail to extract any information, and the failures can be revealed by the MDR to statistically evaluate the corruptions in recovered ECG signals due to noise distortion.
The result of MDR is shown as the cumulative distribution function (CDF) in Figure LABEL:sub@fig:mdr_cdf with the median MDR as $1.7\%$, $0.13\%$ and $0.13\%$ for MMECG, radarODE and radarODE-MTL respectively, and $\Delta m\%$ across $91$ trials are both $14\%$. The reason for the similar performance of two ODE-based methods is that the misaligned ECG pieces with small deviations ($<150$ms) in radarODE will not be identified as ‘missed detected’, and hence the CDFs of MDR share a similar pattern and trend in Figure LABEL:sub@fig:mdr_cdf.
@@ IV-C2 Corrupt ECG Reconstruction
IV-C3 Coarse Cardiac Feature Reconstruction
All three frameworks evaluated in this paper are designed for fine-grained cardiac features reconstruction and should perform well on the coarse cardiac feature (i.e., heart rate (HR) monitoring). The result in Figure LABEL:sub@fig:hr_cdf coincides with the expectation with median HR error as $0.6$, $0.3$ and $0.3$ beats/min respectively, and $\Delta m\%$ for the ODE-based methods are $54\%$ and $59\%$. It is notable in Figure LABEL:sub@fig:hr_cdf that the performances of ODE-based methods are very similar at the beginning, while the radarODE tends to get more errors when the noise in the raw radar signal affects the R peaks recovery, because the calculation of HR is based on the R peak positions.
IV-C4 Fine-Grained Morphological Feature Reconstruction
The morphological feature is an essential fine-grained feature to describe the general similarity between the recovered and ground truth ECG signals, and the morphological accuracy can be evaluated by RMSE, PCC and $R^{2}$, with RMSE sensitive to the peak deviation, PCC focusing on the similarity of the general shape and $R^{2}$ shows the interpretability of the well-trained neural network. The results are shown in Figure LABEL:sub@fig:rmse_cdfLABEL:sub@fig:ocor_cdf and LABEL:sub@fig:r2_cdf as the CDF of RMSE/PCC/$R^{2}$ across $91$ trials in the dataset, and three frameworks get the median RMSE/PCC as $0.125$mV/$82.1\%$/$0.74$, $0.098$mV/$90.1\%$/$0.81$ and $0.083$mV/$92.7\%$/$0.85$ respectively.
As indicated by $\Delta m\%$, the improvements of RMSE ($28\%,33\%$) are larger than PCC ($18\%,21\%$) for radarODE and radarODE-MTL respectively, because the ODE model embedded in the decoder preserves the main features of ECG even under noises and contributes more on the peaks than on the shapes. In addition, radarODE-MTL further improves the results by aligning the ECG pieces with the predicted anchors, avoiding the misalignment issue in radarODE. Lastly, the resultant improvements in $R^{2}$ ($17\%$ and $19\%$ in Figure LABEL:sub@fig:r2_cdf) indicate that radarODE-MTL could capture more dependency in the domain transformation of cardiac activities due to the induction of ODE model as prior knowledge, but the improvements are less than the other two metrics for morphological assessment, because $R^{2}$ is not sensitive to the outliers and could objectively evaluate the model ability.
IV-C5 Fine-Grained ECG Peaks Reconstruction
@@ IV-D Noise Robustness Test
In this work, $10$ trials (No. $75-84$) are selected for the noise robustness test by adding different types of synthesized noises with certain decibel (dB) only in the test stage, while the training and validation stage will use the original data because adding noises into the training dataset is a data augmentation technique to improve the model performance, causing an unfair comparison in the noise robustness test [32]. In addition, adding noises into the validation stage is equivalent to selecting appropriate models for the scenarios with different SNR levels and cannot prove the noise robustness of the proposed radarODE-MTL.
@@ IV-D1 Constant Noise
The constant noise normally affects the SNR of the signal and could be caused by thermal noise from electronic components or long-range detection [17, 6], e.g., the SNR for the current data collection scenario will decrease to 00dB by increasing the monitoring distance to $5$m. In the literature, low SNR scenarios can be simulated by adding Gaussian noise with different intensities as implemented in [6, 20, 43, 41]. The baseline results for three frameworks are firstly obtained in terms of the RMSE, PCC, $R^{2}$, R-peak error and MDR as shown in Table III, and $\Delta m\%$ is calculated as $0\%$, $7.47\%$ and $10.64\%$ as indicated by the initial points in Figure 9. Then, the Gaussian noises with $6$ to $-3$dB are added into the raw radar signal without retraining the deep-learning framework, and the results are shown in Table III with the trends of performance degradation shown in Figure 9.
@@
SNR RMSE (mV) $\downarrow$ PCC $\uparrow$ $R^{2}$uparrow Peak Error (ms) $\downarrow$ MDR $\downarrow$ $\Delta m\%^{1}\uparrow$ $P$ Value (×$\times 10^{-2}$)
@@
Baseline 0.107 83.75% 0.77 9.45 4.52% 0.00$\pm$0.55 -
${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}6$ dB 0.107 82.60% 0.76 9.76 4.37% -0.28$\pm$1.68 73.85
${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}3$ dB 0.108 82.64% 0.76 9.85 4.84% -4.17$\pm$1.91 0.04
${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}0$ dB 0.109 80.00% 0.74 11.80 4.92% -12.38$\pm$3.18 0.00
$-1$ dB 0.114 78.55% 0.69 12.20 5.32% -18.17$\pm$4.10 0.00
$-2$ dB 0.120 74.32% 0.65 14.64 5.59% -30.53$\pm$3.78 0.00
$-3$ dB 0.127 62.45% 0.54 21.28 6.40% -63.81$\pm$2.15 0.00
@@
Baseline 0.091 83.53% 0.79 9.08 4.03% 0.00$\pm$0.39 -
${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}6$ dB 0.093 83.30% 0.78 9.12 4.36% -3.29$\pm$1.71 0.11
${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}3$ dB 0.095 83.01% 0.76 9.01 4.70% -6.26$\pm$1.43 0.00
${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}0$ dB 0.101 82.21% 0.69 9.89 5.86% -20.91$\pm$2.09 0.00
$-1$ dB 0.116 79.66% 0.63 11.90 5.36% -27.17$\pm$2.92 0.00
$-2$ dB 0.157 70.87% 0.58 13.95 6.19% -48.44$\pm$3.53 0.00
$-3$ dB - - - - - Failed2 -
@@
Baseline 0.089 85.03% 0.81 8.22 4.08% 0.00$\pm$1.24 -
${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}6$ dB 0.088 85.31% 0.82 8.18 4.20% -0.52$\pm$0.72 19.81
${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}3$ dB 0.089 84.29% 0.80 8.31 4.27% -2.15$\pm$1.55 0.78
${\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}-}0$ dB 0.091 83.77% 0.79 8.03 4.76% -5.47$\pm$2.59 0.05
$-1$ dB 0.093 84.01% 0.78 8.10 5.10% -8.89$\pm$1.21 0.00
$-2$ dB 0.093 84.51% 0.79 8.02 5.45% -11.22$\pm$1.43 0.00
$-3$ dB 0.094 84.96% 0.78 8.19 6.02% -16.77$\pm$1.60 0.00
@@
2. The ECG recovery fails if PCC$<60\%$, according to the empirical observation of the morphological ECG features.
A general observation of Table III is that all the frameworks perform well before 00dB with a similar degradation rate as in Figure LABEL:sub@fig:constant. Then, radarODE-MTL could still provide reasonable results with mild degradation after 00dB because the MTL paradigm split the ECG reconstruction task into several sub-tasks, and each task can either be constrained by prior knowledge or leverage the information from context data with less pollution. In contrast, radarODE could generate high-fidelity ECG pieces as claimed in [5] and gets the second best baseline result in Table III, but the design of PPI estimation stage does not consider the noise robustness. Therefore, the performance is heavily dropped to the worst in Figure LABEL:sub@fig:constant because of the bad results of Peak Error as shown in Table III. Lastly, the MMECG considers the ECG recovery as an arbitrary domain transformation problem without any constraints in the network design, and the performance also heavily degrades in Figure LABEL:sub@fig:constant because only meaningless results will be generated as shown previously in Figure LABEL:sub@fig:mmecg_abrupt.
@@
Duration RMSE (mV) $\downarrow$ PCC $\uparrow$ $R^{2}$uparrow Peak Error (ms) $\downarrow$ MDR $\downarrow$ $\Delta m\%^{1}\uparrow$    RMSE (mV) $\downarrow$ PCC $\uparrow$ $R^{2}$uparrow Peak Error (ms) $\downarrow$ MDR $\downarrow$ $\Delta m\%\uparrow$
@@ IV-D2 Abrupt Noise
In this part, the Gaussian noises with different intensities (00 and $-9$dB) are used to simulate mild body movement (e.g., during talking or writing) and extensive body movement (e.g., during torso movement) as suggested in the literature [44]. In practice, the body movements have orders of magnitude larger than cardiac activities to ruin the cardiac activities, and the ability of radarODE-MTL to recover ECG signal during RBM comes from the contextual information provided by previous cardiac cycles without RBM noise. Only $20\%$ of the segments randomly selected from one trial are doped, and the duration of noise varies from $1$ to $3$ sec.
For mild body movement, the experimental results are shown in Table IV with the changes of $\Delta m\%$ shown in Figure LABEL:sub@fig:abrupt. Firstly, it is evident that the impact of $1$-sec abrupt noise is limited for all the frameworks, and the results for ODE-based methods are almost equivalent to the baselines. Secondly, $2$-sec noise starts to have a noticeable impact on MMECG, while the ODE-based methods could preserve the performance on the morphological features (RMSE/PCC/$R^{2}$) with small degradation on the Peak Error and MDR. Lastly, $3$-sec noise has distorted $3/4$ of the input radar segment, and the performances of MMECG and radarODE drop obviously as shown in Figure LABEL:sub@fig:abrupt, while radarODE-MTL only loses some points on $\text{MDR}=5.12\%$ as shown in Table IV.
In comparison, the extensive body movements with $1$ and $2$ sec have similar impacts with mild ones on ODE-based methods, because the ODE decoder could preserve the ECG shape even under strong noises, whereas the segments affected by noise cannot contribute to the recovery for MMECG as evident by the significant drop of PCC (from $84.05\%$ to $79.01\%$) as shown in Table IV. In addition, the $3$-sec noise destroys the ECG recovery for MMECG and radarODE with a significant degradation as shown in Figure LABEL:sub@fig:abrupt, whereas the radarODE-MTL only sacrifices certain RMSE and peak accuracy with the overall degradation dropping slightly from $-9.72\%$ to $-10.60\%$ as shown in Table IV.
Statistical Analysis for Noise Robustness Test: The same T-test is also implemented for the noise robustness test as shown in Table III and IV. For constant noise, $6$dB noise does not have a significant impact on the performance of MMECG and radarODE-MTL with $P>0.05$, while the performance of radarODE degrades because of the PPI error accumulation [5]. In addition, all the $P$ values for the experiments of abrupt noise are less than $0.05$ and are not listed in Table IV. Lastly, the impact of the noise level or duration is statistically significant, and the proposed radarODE-MTL could significantly improve the noise robustness as shown in Table III and IV, because the CIs of all $\Delta m\%$ have no overlapping with MMECG or radarODE under the same noise.
@@
IV-E Complexity Analysis and Comparison
Table V presents a detailed complexity comparison of three frameworks considering the parameter count (Params.), floating point operations (FLOPs), multiply-accumulate operations (MACs), and training time per epoch. The parameter count reflects the total number of parameters in each model, and FLOPs and MACs quantify the computational costs [11]. As shown in Table V, the complexities of ODE-based methods are higher than that of MMECG due to the different input data types, with radarODE and radarODE-MTL using spectrogram input and MMECG processing 1D radar signals. In addition, the majority of the parameters ($59\%$) and FLOPs ($95\%$) for radarODE-MTL are for the backbone stage for spectrogram processing, and an important future work is to squeeze the backbone size with reduced input spectrograms.
Compared with the gaps in model size, the training times per epoch for three frameworks are closer, because the MMECG is trained on arbitrary radar/ECG segments with a step length of $0.15$ sec [2], while the ODE-based frameworks are based on single cardiac cycles. In this case, the MMECG needs to traverse $48$k samples while radarODE-MTL only has $19$k samples, indicating that many samples for MMECG training are homogeneous and cannot contribute to dataset diversity and may increase the risk of overfitting.
TABLE V: Complexity Comparison Across Deep Learning Frameworks
Framework Params. (M) FLOPs (G) MACs (G) Time/Epoch (min)
MMECG [2] $0.67$ $0.59$ $0.30$ $3.25$
radarODE [5] $6.04$ $2.45$ $1.23$ $4.51$
radarODE- MTL Backbone $4.81$ $2.37$ $1.18$
Encoder $0.72$ $0.05$ $0.03$
Decoder $2.59$ $0.07$ $0.03$
All $8.12$ $2.50$ $1.23$ $4.85$
IV-F Discussions and Future work
The proposed radarODE-MTL framework has demonstrated superior performance compared to previous approaches in generating reliable ECG signals under noisy conditions. However, potential limitations will be discussed in this subsection to motivate future enhancements in radar-based ECG recovery for practical, real-world scenarios and applications.
IV-F1 Clock Synchronization During Data Collection
In Figure 6, it is obvious that all the ECG R-peaks lag the radar signal peaks in the dataset [2], while the actual ECG signal should lead radar signal for several milliseconds due to the electromyographic activation time (EMAT) [45, 46]. The misalignment is blamed for poor synchronization between the devices for collecting radar and ECG signal, while such clock synchronization is commonly neglected because the essential features of ECG (e.g., shape, peak-to-peak interval) will not be affected by EMAT [2, 47, 41]. In future work, strict clock synchronization should be ensured during data collection to provide faithful radar-ECG pairs for the diagnosis of more diseases with irregular EMAT (e.g., heart failure syndromes and paroxysmal atrial fibrillation) [45, 46].
IV-F2 Robustness During Continuous Large-scale Body Movement
The noise robustness test in Section IV-D shows the better performance of radarODE-MTL compared with other frameworks, because radarODE-MTL could leverage the information from adjacent clean cardiac cycles without noise distortion. However, the recovery may still have poor quality (i.e., bad MDR in Figure 9b) due to the continuous large-scale body movement. For example, if the majority of the input radar signal is contaminated by strong noise without containing any clean cardiac cycle, the deep learning model may not extract any useful information for ECG recovery. In future work, advanced signal processing algorithms are necessary to be developed to ensure a high SNR signal even under continuous large-scale body movement to enable radar-based cardiac monitoring in a general scenario (e.g., walking subjects).